# SenseNova-MARS: Empowering Multimodal Agentic Reasoning and Search via Reinforcement Learning

### v1

@@ Abstract
While Vision-Language Models (VLMs) can solve complex tasks through agentic reasoning, their capabilities remain largely constrained to text-oriented chain-of-thought or isolated tool invocation. They fail to exhibit the human-like proficiency required to seamlessly interleave dynamic tool manipulation with continuous reasoning, particularly in knowledge-intensive and visually complex scenarios that demand coordinated external tools such as search and image cropping. In this work, we introduce SenseNova-MARS, a novel Multimodal Agentic Reasoning and Search framework that empowers VLMs with interleaved visual reasoning and tool-use capabilities via reinforcement learning (RL). Specifically, SenseNova-MARS dynamically integrates the image search, text search, and image crop tools to tackle fine-grained and knowledge-intensive visual understanding challenges. In the RL stage, we propose the Batch-Normalized Group Sequence Policy Optimization (BN-GSPO) algorithm to improve the training stability and advance the model’s ability to invoke tools and reason effectively. To comprehensively evaluate the agentic VLMs on complex visual tasks, we introduce the HR-MMSearch benchmark, the first search-oriented benchmark composed of high-resolution images with knowledge-intensive and search-driven questions. Experiments demonstrate that SenseNova-MARS achieves state-of-the-art performance on open-source search and fine-grained image understanding benchmarks. Specifically, on search-oriented benchmarks, SenseNova-MARS-8B scores 67.84 on MMSearch and 41.64 on HR-MMSearch, surpassing proprietary models such as Gemini-3-Flash and GPT-5. SenseNova-MARS represents a promising step toward agentic VLMs by providing effective and robust tool-use capabilities. To facilitate further research in this field, we will release all code, models, and datasets.
@@ 1 Introduction
Figure 1: Overall performance of SenseNova-MARS-8B compares to other models across six benchmarks. All other models are evaluated under agentic workflow. SenseNova-MARS-8B demonstrates exceptional performance on the search-oriented benchmarks such as MMSearch [jiang2024mmsearch], HR-MMSearch and FVQA [wu2025mmsearch], surpassing leading proprietary models such as Gemini-3-Flash [gemini3flash] and GPT-5 [openai2025gpt5]. For the high-resolution perception benchmark such as V* Bench [wu2024v] and HR-Bench [wang2025divide], SenseNova-MARS-8B also outperforms existing open-source models, including DeepEyesV2 [hong2025deepeyesv2] and Mini o3 [lai2025mini].
@@
To this end, we propose SenseNova-MARS, a novel Multimodal Agentic Reasoning and Search framework that leverages RL to integrate image search, text search, and image crop tools into a dynamic, multi-turn reasoning process. Specifically, SenseNova-MARS adaptively interacts with the tool sets through reasoning and planning, learning when and how to invoke the search and image crop tools during the iterative reasoning process. We systematically investigate how to build this agentic VLM based on Qwen2.5-VL-7B [bai2025qwen2] and Qwen3-VL-8B [Qwen3-VL], named SenseNova-MARS-7B and SenseNova-MARS-8B, respectively. Our approach focuses on three key aspects: training data construction, a two-stage training pipeline including the cold-start stage and RL, and the design of RL algorithm. We first curate the high-quality cold-start and RL training data through a synthesis pipeline with rigorous quality verification. The dataset spans diverse visual tasks, including knowledge-intensive, high-resolution perception, and search-oriented tasks, which necessitate the use of the search tool, image crop tool, or a combination of both. During the initial cold-start supervised fine-tuning (SFT) stage, the model learns basic tool-usage patterns from a minimal dataset of approximately 3,000 samples. This compact dataset is crucial as it establishes a foundation for the subsequent RL stage, particularly in learning to use the tools effectively. To advance the agentic VLM’s capabilities in multi-tool collaboration and reasoning during the RL stage, we propose Batch-Normalized Group Sequence Policy Optimization (BN-GSPO), an extension of the standard GSPO algorithm. Experiments indicate that BN-GSPO can improve training stability for multi-tool rollout trajectories and yield significant performance gains.
To further evaluate visual search agents for complex scenes, we introduce HR-MMSearch, the first search-oriented benchmark composed of high-resolution images with knowledge-intensive and search-driven questions. SenseNova-MARS demonstrates superior performance on HR-MMSearch, far beyond the search-only or crop-only models, and is comparable with proprietary models like Gemini-2.5-Flash [comanici2025gemini], Gemini-3-Flash [gemini3flash], and GPT-5 [openai2025gpt5]. Furthermore, SenseNova-MARS-8B achieves state-of-the-art (SOTA) results across diverse open-source search-oriented and perception benchmarks as shown in Fig. 1. It outperforms MMSearch-R1 [wu2025mmsearch] and DeepEyesV2 [hong2025deepeyesv2] on search-oriented benchmarks and attains a leading score of 67.84 in MMSearch [jiang2024mmsearch] and 41.64 in HR-MMSearch. For the high-resolution perception benchmark, SenseNova-MARS-8B achieves 92.2 on V* Bench [wu2024v] and 83.1 on HR-Bench 4K [wang2025divide], exceeding specialized models like DeepEyes [zheng2025deepeyes], Pixel-Reasoner [su2025pixel] and Mini o3 [lai2025mini].
@@ 3.4 HR-MMSearch Benchmark Construction
GPT-5 [openai2025gpt5] 50.24 35.09 22.62 54.39 61.70 54.15 44.39 79.33
Gemini-3-Flash [gemini3flash] 53.68 57.31 21.97 56.50 63.57 54.85 38.90 82.67
@@
GPT-5 [openai2025gpt5] 60.12 52.63 38.36 62.61 70.58 55.95 56.02 84.67
Gemini-3-Flash [gemini3flash] 61.26 62.57 41.64 64.89 67.92 61.10 48.06 82.67
@@ 4.1 Implementation Details
Model and Training. SenseNova-MARS-7B, which is based on Qwen2.5-VL-7B-Instruct [bai2025qwen2], is trained using the two-stage pipeline. This process involves SFT followed by RL, which we implement using the LLaMA-Factory [zheng2024llamafactory] and veRL [sheng2025hybridflow] frameworks, respectively. During the SFT stage, we fine-tune only the language model while keeping the vision encoder and multi-modal projector frozen. The training uses a learning rate of $1\times 10^{-5}$ for 3 epochs. The SFT stage bootstraps the model’s ability to follow the interaction protocol and utilize the multi-toolset. The subsequent RL stage then teaches the model to use these tools more effectively and efficiently. For RL stage, we use a global batch size of 128, a learning rate of $1\times 10^{-6}$, and a KL coefficient $\beta$ of $1\times 10^{-4}$. Additionally, to encourage better exploration, we follow DAPO [yu2025dapo] by using the Clip-Higher strategy with $\epsilon_{\text{low}}=0.2$ and $\epsilon_{\text{high}}=0.28$. During this RL phase, a single training trajectory allows the agent to interact for a maximum of $T=10$ turns. This means the agent can iteratively use tools and refine its plan for up to 10 steps before producing a final answer for that trajectory. In each turn, the agent can generate up to $8,192$ tokens, with a cumulative limit of $32,768$ tokens for the entire trajectory. Compared with SenseNova-MARS-7B, SenseNova-MARS-8B is developed based on Qwen3-VL-8B-Instruct [Qwen3-VL], and only the RL stage is employed, leveraging the strong tool-usage capability of the base model. Detailed training hyperparameters are available in the Appendix.
@@
Baselines. We evaluate our model against several strong baselines. These include proprietary models, such as GPT-4o [hurst2024gpt], GPT-5 [openai2025gpt5], Gemini-2.5-Flash [comanici2025gemini] and Gemini-3-Flash [gemini3flash], as well as open-source agentic models like MMSearch-R1 [wu2025mmsearch], DeepMMSearch-R1 [narayan2025deepmmsearch] and DeepEyesV2 [hong2025deepeyesv2]. All models are tested under the following two workflow settings:
@@ 4.2 Main Results
Search-oriented Evaluation. As shown in Tab. 1, SenseNova-MARS-8B establishes a new SOTA among open-source agentic models under 8B parameters and closed-source proprietary models. Specifically, SenseNova-MARS-8B attains an average performance gain of 12.68 points compared with Qwen3-VL-8B. Additionally, SenseNova-MARS outperforms search-orientSearch-oriented Evaluationed agentic models such as MMSearch-R1 [wu2025mmsearch], DeepMMSearch-R1 [narayan2025deepmmsearch], and DeepEyesV2 [hong2025deepeyesv2], surpassing MMSearch-R1 by an average of 11.71 points. Compared with proprietary closed-source models, SenseNova-MARS-8B surpasses GPT-5 [openai2025gpt5], Gemini-2.5-Flash [comanici2025gemini], and Gemini-3-Flash [gemini3flash] by a significant margin, outperforming Gemini-3-Flash by an average of 2.94 points. These results highlight the efficacy of our unified agentic search and visual reasoning RL framework.
Fine-grained Visual Understanding. As shown in Tab. 2, SenseNova-MARS-8B demonstrates superior fine-grained perception capabilities on high-resolution benchmarks, validating the effectiveness of our approach in detailed, pixel-space analysis. SenseNova-MARS-8B achieves leading scores of 92.2 on V* Bench [wu2024v], 83.1 on HR-Bench 4k [wang2025divide], 78.4 on HR-Bench-8k [wang2025divide], and 67.9 on MME-RealWorld [zhang2024mme], outperforming all existing tool-based models such as Pixel Reasoner [su2025pixel], DeepEyes [zheng2025deepeyes], and Mini o3 [lai2025mini]. Compared with Qwen3-VL-8B, SenseNova-MARS-8B attains an average performance gain of 4.9 points, demonstrating the effectiveness of our proposed RL algorithm.

### v2

@@ Abstract
While Vision-Language Models (VLMs) can solve complex tasks through agentic reasoning, their capabilities remain largely constrained to text-oriented chain-of-thought or isolated tool invocation. They fail to exhibit the human-like proficiency required to seamlessly interleave dynamic tool manipulation with continuous reasoning, particularly in knowledge-intensive and visually complex scenarios that demand coordinated external tools such as search and image cropping. In this work, we introduce SenseNova-MARS, a novel Multimodal Agentic Reasoning and Search framework that empowers VLMs with interleaved visual reasoning and tool-use capabilities via reinforcement learning (RL). Specifically, SenseNova-MARS dynamically integrates the image search, text search, and image crop tools to tackle fine-grained and knowledge-intensive visual understanding challenges. In the RL stage, we propose the Batch-Normalized Group Sequence Policy Optimization (BN-GSPO) algorithm to improve the training stability and advance the model’s ability to invoke tools and reason effectively. To comprehensively evaluate the agentic VLMs on complex visual tasks, we introduce the HR-MMSearch benchmark, the first search-oriented benchmark composed of high-resolution images with knowledge-intensive and search-driven questions. Experiments demonstrate that SenseNova-MARS achieves state-of-the-art performance on open-source search and fine-grained image understanding benchmarks. Specifically, on search-oriented benchmarks, SenseNova-MARS-32B scores 74.3 on MMSearch and 54.4 on HR-MMSearch, surpassing proprietary models such as Gemini-3-Pro and GPT-5.2. SenseNova-MARS represents a promising step toward agentic VLMs by providing effective and robust tool-use capabilities. To facilitate further research in this field, we will release all code, models, and datasets.
@@ 1 Introduction
Figure 1: Overall performance of SenseNova-MARS-32B compares to other models across six benchmarks. SenseNova-MARS-32B demonstrates exceptional performance on the search-oriented benchmarks such as MMSearch [jiang2024mmsearch], HR-MMSearch and FVQA [wu2025mmsearch], surpassing leading proprietary models such as Gemini-3-Pro [gemini3pro] and GPT-5.2 [openai2025gpt5_2]. For the high-resolution perception benchmark such as V* Bench [wu2024v] and HR-Bench [wang2025divide], SenseNova-MARS-32B can outperform top-tier models such as Qwen3-VL-235B-A22B [Qwen3-VL].
@@
To this end, we propose SenseNova-MARS, a novel Multimodal Agentic Reasoning and Search framework that leverages RL to integrate image search, text search, and image crop tools into a dynamic, multi-turn reasoning process. Specifically, SenseNova-MARS adaptively interacts with the tool sets through reasoning and planning, learning when and how to invoke the search and image crop tools during the iterative reasoning process. We systematically investigate how to build this agentic VLM based on Qwen2.5-VL [bai2025qwen2], and Qwen3-VL [Qwen3-VL], introducing SenseNova-MARS in three scales including SenseNova-MARS-7B, SenseNova-MARS-8B and SenseNova-MARS-32B. Our approach focuses on three key aspects: training data construction, a two-stage training pipeline including the cold-start stage and RL, and the design of RL algorithm. We first curate the high-quality cold-start and RL training data through a synthesis pipeline with rigorous quality verification. The dataset spans diverse visual tasks, including knowledge-intensive, high-resolution perception, and search-oriented tasks, which necessitate the use of the search tool, image crop tool, or a combination of both. During the initial cold-start supervised fine-tuning (SFT) stage, the model learns basic tool-usage patterns from a minimal dataset of approximately 3,000 samples. This compact dataset is crucial as it establishes a foundation for the subsequent RL stage, particularly in learning to use the tools effectively. To advance the agentic VLM’s capabilities in multi-tool collaboration and reasoning during the RL stage, we propose Batch-Normalized Group Sequence Policy Optimization (BN-GSPO), an extension of the standard GSPO algorithm. Experiments indicate that BN-GSPO can improve training stability for multi-tool rollout trajectories and yield significant performance gains.
To further evaluate visual search agents for complex scenes, we introduce HR-MMSearch, the first search-oriented benchmark composed of high-resolution images with knowledge-intensive and search-driven questions. SenseNova-MARS demonstrates superior performance on HR-MMSearch, far beyond the search-only or crop-only models, and is comparable with proprietary models like Gemini-3-Pro [gemini3pro], and GPT-5.2 [openai2025gpt5_2]. Furthermore, SenseNova-MARS-32B achieves state-of-the-art (SOTA) results across diverse open-source search-oriented and perception benchmarks as shown in Fig. 1, attaining a leading score of 74.3 in MMSearch [jiang2024mmsearch] and 54.4 in HR-MMSearch. For the high-resolution perception benchmark, SenseNova-MARS-32B achieves 94.2 on V* Bench [wu2024v] and 90.2 on HR-Bench 4K [wang2025divide], exceeding top-tier models like Qwen3-VL-235B-A22B [Qwen3-VL] and Skywork-R1V4-30B [zhang2025skywork]. Similarly, our smaller variant, SenseNova-MARS-8B, is superior to Gemini-3-Flash [gemini3flash] and GPT-5 [openai2025gpt5], and achieves SOTA among similarly-sized open-source models such as MMSearch-R1 [wu2025mmsearch] and DeepEyesV2 [hong2025deepeyesv2].
@@ 3.4 HR-MMSearch Benchmark Construction
GPT-5 [openai2025gpt5] 50.24 35.09 22.62 54.39 54.15 61.70 44.39 79.33
GPT-5.2 [openai2025gpt5] 50.92 43.27 24.92 50.94 50.40 59.92 47.00 80.00
Gemini-3-Flash [gemini3flash] 53.68 57.31 21.97 56.50 54.85 63.57 38.90 82.67
Gemini-3-Pro [gemini3pro] 55.87 62.57 26.89 59.22 56.30 64.07 40.06 82.00
@@
GPT-5 [openai2025gpt5] 60.12 52.63 38.36 62.61 55.95 70.58 56.02 84.67
Gemini-3-Flash [gemini3flash] 61.26 62.57 41.64 64.89 61.10 67.92 48.06 82.67
GPT-5.2 [openai2025gpt5] 67.64 66.08 48.20 68.78 65.55 78.18 65.99 80.67
Gemini-3-Pro [gemini3pro] 69.06 74.27 48.52 72.61 66.45 75.91 59.69 86.00
@@ 4.1 Implementation Details
Model and Training. SenseNova-MARS-7B, which is based on Qwen2.5-VL-7B-Instruct [bai2025qwen2], is trained using the two-stage pipeline. This process involves SFT followed by RL, which we implement using the LLaMA-Factory [zheng2024llamafactory] and veRL [sheng2025hybridflow] frameworks, respectively. During the SFT stage, we fine-tune only the language model while keeping the vision encoder and multi-modal projector frozen. The training uses a learning rate of $1\times 10^{-5}$ for 3 epochs. The SFT stage bootstraps the model’s ability to follow the interaction protocol and utilize the multi-toolset. The subsequent RL stage then teaches the model to use these tools more effectively and efficiently. For RL stage, we use a global batch size of 128, a learning rate of $1\times 10^{-6}$, and a KL coefficient $\beta$ of $1\times 10^{-4}$. Additionally, to encourage better exploration, we follow DAPO [yu2025dapo] by using the Clip-Higher strategy with $\epsilon_{\text{low}}=0.2$ and $\epsilon_{\text{high}}=0.28$. During this RL phase, a single training trajectory allows the agent to interact for a maximum of $T=10$ turns. This means the agent can iteratively use tools and refine its plan for up to 10 steps before producing a final answer for that trajectory. In each turn, the agent can generate up to $8,192$ tokens, with a cumulative limit of $32,768$ tokens for the entire trajectory. Compared with SenseNova-MARS-7B, SenseNova-MARS-8B and SenseNova-MARS-32B are developed based on Qwen3-VL-8B-Instruct and Qwen3-VL-32B-Instruct [Qwen3-VL], respectively, and only the RL stage is employed, leveraging the strong tool-usage capability of the base model. Detailed training hyperparameters are available in the Appendix.
@@
Baselines. We evaluate our model against several strong baselines. These include proprietary models, such as GPT-5 [openai2025gpt5], GPT-5.2 [openai2025gpt5_2], Gemini-3-Flash [gemini3flash] and Gemini-3-Pro [gemini3pro], as well as open-source agentic models like MMSearch-R1 [wu2025mmsearch], DeepMMSearch-R1 [narayan2025deepmmsearch] and DeepEyesV2 [hong2025deepeyesv2]. All models are tested under the following two workflow settings:
@@
Qwen3-VL-32B-Instruct [Qwen3-VL] 91.1 84.6 81.6 - -
Qwen3-VL-235B-A22B-Instruct [Qwen3-VL] 93.7 85.4 82.4 - -
@@ 4.2 Main Results
Search-oriented Evaluation. As shown in Tab. 1 and Fig. 6, SenseNova-MARS-32B establishes a new SOTA among open-source agentic models and closed-source proprietary models. SenseNova-MARS-32B achieves an average score of 69.74 across seven benchmarks, surpassing leading proprietary models such as Gemini-3-Pro [gemini3pro] and GPT-5.2 [openai2025gpt5_2]. Notably, on the HR-MMSearch benchmark, SenseNova-MARS-32B scores 54.43, outperforming Gemini-3-Pro and GPT-5.2 by margins of 5.91 and 6.23 points, respectively. In the small-scale category, SenseNova-MARS-8B also demonstrates strong performance, attaining an average gain of 12.68 points over Qwen3-VL-8B. Additionally, SenseNova-MARS outperforms search-oriented agentic models such as MMSearch-R1 [wu2025mmsearch], DeepMMSearch-R1 [narayan2025deepmmsearch], and DeepEyesV2 [hong2025deepeyesv2], surpassing MMSearch-R1 by an average of 11.71 points. Compared with proprietary closed-source models, SenseNova-MARS-8B surpasses GPT-5 [openai2025gpt5], Gemini-2.5-Flash [comanici2025gemini], and Gemini-3-Flash [gemini3flash] by a significant margin, outperforming Gemini-3-Flash by an average of 2.94 points. These results highlight the efficacy of our unified agentic search and visual reasoning RL framework.
Fine-grained Visual Understanding. As shown in Tab. 2 and Fig. 6, SenseNova-MARS demonstrates superior fine-grained perception capabilities on high-resolution benchmarks, validating the effectiveness of our approach in detailed, pixel-space analysis. SenseNova-MARS-32B scores 94.2 on V* Bench [wu2024v] and 90.2 on HR-Bench 4k [wang2025divide], surpassing top-tier models such as Qwen3-VL-235B-A22B-Instruct [Qwen3-VL] and Gemini-2.5-Pro [comanici2025gemini]. SenseNova-MARS-8B achieves leading scores of 92.2 on V* Bench [wu2024v], 83.1 on HR-Bench 4k [wang2025divide], 78.4 on HR-Bench-8k [wang2025divide], and 67.9 on MME-RealWorld [zhang2024mme], outperforming all existing tool-based models such as Pixel Reasoner [su2025pixel], DeepEyes [zheng2025deepeyes], and Mini o3 [lai2025mini]. Compared with Qwen3-VL-8B, SenseNova-MARS-8B attains an average performance gain of 4.9 points, demonstrating the effectiveness of our proposed RL algorithm.
Figure 6: Overall performance of SenseNova-MARS-32B and SenseNova-MARS-8B compares to other models.